OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3. y − 6 = −3x − 3 y − 6 = −3(x + 1) y = −3x + 3 3x + y = 3

OpenStudy (anonymous):

2nd one

OpenStudy (anonymous):

ight n i need yall to explain to me

OpenStudy (anonymous):

point-slope form = y-y1= m(x-x1)

hartnn (hartnn):

the slope point form is (y-y1)=m(x-x1) m= slope, (x1,y1) is point

OpenStudy (anonymous):

o second then right??

hartnn (hartnn):

yup

OpenStudy (anonymous):

okay thxs n keep helping me plzz